# Activation of entanglement in generalized entanglement swapping

### v1

@@ I INTRODUCTION
In operational terms, an operator is referred to as an inseparable operator if its Choi matrix is entangled while separable operators are operators which take separable states to separable states and its Choi matrix is a separable density matrix rain1 ; rain2 ; rain3 . Quantum measurements are a special class of operators. The most general form of measurement is called a Positive operator valued measure (POVM) um1 ; unsharp_measurement ; povm . A POVM, $\mathcal{M}$ is defined as a set of positive operators $\Pi_{n}$ s.t $\sum_{n}{\Pi_{n}}=\mathbb{I}$ where $\mathbb{I}$ is the identity operator on the corresponding Hilbert space. Measurements are characterized by various definitions, in papers terhal ; vincenzo ; yokoyama ; hamamura ; sperling ; duan ; vertesi ; rabelo ; adam ; wu ; virmani ; lobet , it has been defined that a POVM, =$\mathcal{M}=\{\Pi_{n}\}$, ($n^{\text{th}}$) POVM element $\Pi_{n|AB}$ is entangled if the corresponding operator defined as $\frac{\Pi_{n|AB}^{T_{B}}}{\mbox{Tr}(\Pi_{n|AB})}$ fails to be positive where $T_{B}$ represent the partial transpose with respect to the subsystem B. The POVM element is said to be unentangled if the normalized form $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$ is a separable quantum mixed state. A POVM measurement is called an entangled measurement if at least one of the POVM elements is entangled. Similarly, A POVM measurement is called unentangled measurement if each of the POVM elements is unentangled. Here, the entanglement of the measurement is characterized as an entanglement of the detection devices or entanglement of the measurement. The entire measurement set is classified into different categories based on this definition vertesi ; adam . In vertesi ; adam , the authors have classified the set of measurements into different categories such as entangled measurement, unentangled measurement, LOCC measurement, and classical measurement where classical and LOCC measurements do not have any capacity to produce the entangled state from a separable state. Note that these measurements have the inclusion relation: General measurement $\supset$ unentangled measurement $\supset$ LOCC measurement $\supset$ classical measurement, where general measurement corresponds to the set of all quantum measurements, including the entangled measurements vertesi ; adam . LOCC and classical measurements belong to the class of separable operations in an operational sense. It is important to note that, according to this definition, not every unentangled measurement is considered a separable operation, but the converse is true.
@@
The standard entanglement swapping procedure has been generalized to work with non-maximally entangled states, and projective measurements have found important applications in quantum networks and nonlocality-related problems perseguers1 ; perseguers2 ; wojcik ; klobus ; gour . In terhal ; vincenzo , it was shown that entangled measurements are necessary and sufficient to obtain an entangled state between Alice and Charlie when the initial states are Bell pairs, and measurements are generalized measurements called POVM.
@@
In this work, we re-evaluate the necessity and sufficiency conditions for measurement operators to be entangled when initial states are maximally entangled and measurements are generalized measurements. We start with two Bell pairs: (1, 2) shared between Alice and Bob, and (3, 4) shared between Bob and Charlie. A quantum measurement, denoted by positive operator-valued measure (POVM), is applied to qubits (2, 3), leading to the creation of a shared state in (1, 4) between the spatially separated observers. We demonstrate that while an entangled measurement is sufficient, it is not necessary for the measurement operators to be entangled to establish an entangled state between A and C. By employing a sequential approach where Bob conducts an initial measurement, and if that does not establish the entanglement, followed by another measurement, it may possible to achieve entanglement. This sequential strategy, combined with appropriate post-processing after the first measurement, enables the establishment of entanglement between A and C. We identify the specific criteria for different measurement operators that enable the possibility of performing a second measurement to establish entanglement. In cases where the first measurement fails to establish entanglement, we delineate the protocol accordingly. We demonstrate that after the first measurement, the $14|23$ bipartition entanglement of the total system state should be between 0 and 1, and the rank of the measurement should be greater than one to enable a disturbance of the (1,4) state by the second measurement. Additionally, we show that a zero $14|23$ bipartition entanglement resulting from a measurement that is an inseparable operator is not feasible. Our demonstration showcases the feasibility of generating entanglement between A and C for different measurement operators. Furthermore, an illustrative example underscores how successive measurements yield more substantial entanglement compared to a single measurement. Specifically, when employing one unentangled measurement incapable of establishing entanglement, a second measurement involving purely separable operation becomes effective. Our approach is not limited to bipartite qubit states but can also be generalized to higher-dimensional maximally entangled states.
@@ II GENERALIZED PROTOCOL
a pure state which is separable in $12|34$ bipartition. Now Bob’s first measurement $\mathcal{M}_{1}$ can be specified by a collection of two-qubit positive-semidefinite operators $\Pi_{n}$ satisfying $\sum_{n=1}^{K}\Pi_{n}=\mathbb{I}_{4}$ where each POVM element $\Pi_{n}$ can be decomposed in the form
@@
where $|\phi_{n\alpha}\rangle$, $\alpha=1,2,3,4$ form a complete orthonormal basis on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $\pi_{n\alpha}\geq 0$ for $\alpha=1,2,3,4$. If Bob performs a joint two-qubit generalized quantum measurement $\mathcal{M}_{1}$ on ($2,3$) of the $|\Phi\rangle$ state then the post-measurement state of the $n$-th outcome ($n\in\{1,...,K\}$) is
@@
where the complex conjugation is in the computation basis. Note that after the measurement, the total system state Eq.5 remains pure.
@@
$\rho_{14|n}$ $=$ $\frac{1}{4p_{n}}\sum_{\alpha=1}^{4}\pi_{n\alpha}|\phi_{n\alpha}^{*}\rangle\langle\phi_{n\alpha}^{*}|$ (6)
@@
where $\Pi_{n}^{*}=\sum_{\alpha=1}^{4}\pi_{n\alpha}|\phi_{n\alpha}^{*}\rangle\langle\phi_{n\alpha}^{*}|$. From Eq.6, we see that the state of the pair (1, 4) is completely determined by Bob’s measurement.
The density of state of the above Eq.7 is
$\rho_{1234|n}$ $=$ $|\Phi_{n}\rangle\langle\Phi_{n}|$ (12)
From the orthogonality condition of the orthonormal basis, we get
$=$ 00 0
which reduce to
$\sum_{i,j=0}^{1}a^{*}_{n\beta|ij}a_{n\alpha|ij}=0$ (14)
$=$ $\frac{1}{\mbox{Tr}(\Pi_{n})}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j}a^{\ast}_{n\beta|k^{\prime}l}|ik\rangle\langle i^{\prime}k^{\prime}|_{12}$ (20)
From the above Eq.6, Eq.17, and Eq.24 one can easily compute the post-measurement states of $n$-th outcome between different pairs (1,4), (1,2), and (3,4) respectively after the $\mathcal{M}_{1}$ POVM measurement.
Equation 6 illustrates the direct relationship between the post-measurement state of the (1,4) pair and the POVM measurement element. Entanglement at the (1,4) pair is a necessary and sufficient condition, as demonstrated in previous works terhal ; vincenzo ; yokoyama . However, we show that while an entangled POVM is sufficient to obtain an entangled state at (1,4), it is not necessary.
Our demonstration reveals that through a sequential approach, where Bob performs an initial measurement $\mathcal{M}_{1}$ and, if necessary, follows it with another measurement $\mathcal{M}_{2}$, entanglement can be achieved. This sequential strategy, coupled with appropriate post-processing after the first measurement, facilitates the establishment of entanglement between parties A and C. Notably, the second measurement $\mathcal{M}_{2}$ does not need to be an inseparable operation; a measurement that is separable operation can also achieve the accomplish the task.
@@
We consider the second measurement $\mathcal{M}_{2}$ is also two qubits generalized measurement (POVM) which can be specified by a collection of two qubits semi-definite operators $E_{m}$ satisfying $\sum_{m=1}^{L}E_{m}=\mathbb{I}_{4}$ where each POVM element $E_{m}$ can be decomposed in the previous form as
@@
Now using Eq.5, and Eq.25, and $|\Phi_{nm}\rangle$ (Eq.26) can be written as
@@
where $|\phi_{n\alpha(\beta)}\rangle$, $\alpha(\beta)=1,2,3,4$ are orthonormal basis of the $n$-th element ($\Pi_{n}$) of the first measurement $\mathcal{M}_{1}$.
@@
Next, to analyze the bipartition entanglement of post-measurement states for each outcome, we consider I-concurrence rungta for higher-dimensional pure quantum states. For defining the bipartition entanglement $A|B$, we employ the definition of I-concurrence as follows:
@@
Proof: Suppose $\mathcal{M}_{1}$ is a projective or rank-1 POVM, implying each element $\Pi_{n}$ of $\mathcal{M}_{1}$ measurement is rank-1, such that for $\alpha=1$, $\Pi_{n}=\pi_{n1}|\phi_{n1}\rangle\langle\phi_{n1}|$. Then, according to Eq.5, $|\Phi_{n}\rangle$ can be expressed as:
@@
Hence, if $\Pi_{n}$ is a projective or rank-1 entangled POVM element operator, the post-measurement state between A and C will be entangled, but no subsequent measurement can alter the entanglement in the (1,4) pair. Similarly, if $\Pi_{n}$ is an unentangled measurement rank-1 or projective element, the post-measurement state between (1,4) will be separable, and no further sequential measurement on the (2,3) pair can entangle the (1,4) pair.
The above proof indicates that if Bob performs the first measurement with a rank greater than one (i.e., $\alpha>1$), there is a possibility to disturb the (1,4) pair using another round of measurement on the (2,3) pair. This is because, from Eq.5, it can be understood that the $14|23$ bipartition may exhibit non-zero entanglement after Bob’s first measurement if its rank is greater than one. Consequently, one might pursue a subsequent measurement on the (2,3) pair to disturb the (1,4) pair.
Lemma-2: After performing the first measurement $\mathcal{M}_{1}$ on the (2,3) pair of the initial state Eq.1, if the bipartition entanglement $\mathcal{C}_{n|14vs23}$ of the four-qubit state $|\Phi_{n}\rangle$ (Eq.5) exhibits non-zero entanglement for a given outcome, i.e., if the $14|23$ bipartition entanglement lies within the range $0<\mathcal{C}_{n|14vs23}(|\Phi_{n}\rangle\langle\Phi_{n}|)\leq 1$, and the rank of the first measurement is greater than one, then the second measurement $\mathcal{M}_{2}$ has the potential to disturb the entanglement of the (1,4) pair
@@
$=$ $\sqrt{\frac{4}{3}\left[1-\frac{\sum_{\alpha=1}^{4}\pi_{n\alpha}^{2}}{\mbox{Tr}(\Pi_{n})^{2}}\right]}$
@@
Initially, in the initial state (Eq.1) $12|34$ bipartition entanglement is zero i.e. $\mathcal{C}_{12vs34}(|\Phi\rangle\langle\Phi|)=0$ and in the $14|23$ bipartition $\mathcal{C}_{14vs23}(|\Phi\rangle\langle\Phi|)=1$. To get (1,4) pair entangled we need a measurement $\mathcal{M}_{1}$ that can produce non-zero entanglement in $12|34$ bipartition. So, until and unless Bob’s first measurement $\mathcal{M}_{1}$ produces entanglement in the $12|34$ bipartition we can not disturb the state between (1,4) pair during the second measurement even if $14|23$ bipartition has non-zero entanglement. Therefore, $\mathcal{C}_{n|12vs34}(|\Phi\rangle\langle\Phi|)>0$ is necessary for any type of measurement. However, this condition alone is not sufficient to entangle the (1,4) pair. There exist unentangled measurements capable of generating non-zero entanglement in the $12|34$ bipartition but unable to establish entanglement between qubits A and C (i.e., in the (1,4) pair). As previously established, if $\Pi_{n}$ is an unentangled element operator, the post-measurement state between the (1,4) pair will also be unentangled.
To get an entangled state between A and C, we have now given the protocols for the whole two qubits measurement of different measurement categories.
@@ II.1 Inseparable operations
In the operational sense, an inseparable operation is defined as follows: Let $\mathcal{M}$ acting on a bipartite system $AB$ appending auxiliary systems $A_{1}$ and $B_{1}$ respectively and letting the channel act on the state $|\phi^{+}\rangle\langle\phi^{+}|_{A_{1}A}\otimes|\phi^{+}\rangle\langle\phi^{+}|_{BB_{1}}$ as
@@
If the entanglement of the state $\rho_{A_{1}ABB_{1}}$ across the bipartition $A_{1}A\ |\ BB_{1}$ is non-zero then the measurement is called an inseparable operation otherwise separable operation cirac .
In our work, we consider two initial states, $|\phi^{+}\rangle$, which are maximally entangled. The entanglement of the four-qubit post-measurement state, $|\Phi_{n}\rangle$ (1), in the $12|34$ bipartition, i.e., $\mathcal{C}_{n|12vs34}(|\Phi_{n}\rangle\langle\Phi_{n}|)$, is the inseparability condition of the $\Pi_{n}$ element of any arbitrary two-qubit generalized measurement from an operational point of view. We find out the exact expression of inseparability criteria cirac for two-qubit arbitrary measurement using the I-concurrence relation rungta which is given by (for details see Appendix A):
@@
$=$ $\sqrt{\frac{4}{3}\left[1-\frac{1}{\mbox{Tr}(\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|ij}a^{*}_{n\eta|kl}a_{n\gamma|k^{\prime}l}\right]}.$

### v2

@@ I INTRODUCTION
In operational terms, an operator is referred to as an inseparable operator if its Choi state is entangled while separable operators are operators which take separable states to separable states and its Choi state is a separable density matrix rain1 ; rain2 ; rain3 . Quantum measurements are a special class of operators. The most general form of measurement is called a Positive Operator-Valued Measure (POVM) um1 ; unsharp_measurement ; povm . A POVM, $\mathcal{M}$ is defined as a set of positive operators $\Pi_{n}$ s.t $\sum_{n}{\Pi_{n}}=\mathbb{I}$ where $\mathbb{I}$ is the identity operator on the corresponding Hilbert space. Measurements are characterized by various definitions. In papers yokoyama ; hamamura ; sperling ; duan ; vertesi ; rabelo ; adam ; wu ; virmani , it has been defined that a POVM element $\Pi_{n}$ is considered entangled if the partial transpose of the corresponding operator, defined as $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$, fails to be positive. The POVM element is said to be unentangled if the normalized form $\frac{\Pi_{n}}{\mbox{Tr}(\Pi_{n})}$ is a separable quantum mixed state. A POVM measurement is called an entangled measurement if at least one of the POVM elements is entangled. Similarly, A POVM measurement is called unentangled measurement if each of the POVM elements is unentangled. Here, the entanglement of the measurement is characterized as an entanglement of the detection devices or POVM entanglement. The entire measurement set is classified into different categories based on this definition vertesi ; adam . In vertesi ; adam , the authors have classified the set of measurements into different categories such as entangled measurement, unentangled measurement, LOCC measurement, and classical measurement where classical and LOCC measurements do not have any capacity to produce the entangled state from a separable state. Note that these measurements have the inclusion relation: General measurement $\supset$ unentangled measurement $\supset$ LOCC measurement $\supset$ classical measurement, where general measurement corresponds to the set of all quantum measurements, including the entangled measurements vertesi ; adam . LOCC and classical measurements belong to the class of separable operations in an operational sense. It is important to note that, according to this definition, not every unentangled measurement is considered a separable operation, but the converse is true.
@@
The standard entanglement swapping procedure has been generalized to work with non-maximally entangled states, and projective measurements have found important applications in quantum networks and nonlocality-related problems perseguers1 ; perseguers2 ; wojcik ; klobus ; gour ; verma . Recent studies have proposed elegant protocols leveraging POVM measurements for entanglement distribution in quantum networks halder1 ; halder2 . The common understanding suggests that entangled measurements are required for establishing entanglement between distant parties yokoyama ; perseguers2 .
@@
In this work, we re-evaluate the necessity and sufficiency conditions for measurement operators to be entangled when initial states are maximally entangled and measurements are generalized measurements. We start with two Bell pairs: (1, 2) shared between Alice and Bob, and (3, 4) shared between Bob and Charlie. A POVM measurement is applied to qubits (2, 3), leading to the creation of a shared state in (1, 4) between the spatially separated observers. We demonstrate that while an entangled measurement is sufficient, it is not necessary for the measurement operators to be entangled to establish an entangled state between A and C. By employing a sequential approach where Bob conducts an initial measurement, and if that does not establish the entanglement, followed by another measurement, it is possible to achieve entanglement. This sequential strategy, combined with appropriate post-processing after the first measurement, enables the establishment of entanglement between A and C. We identify the specific criteria for different measurement operators that enable the possibility of performing a second measurement to establish entanglement. In cases where the first measurement fails to establish entanglement, we delineate the protocol accordingly. We demonstrate that after the first measurement, the $14|23$ bipartition entanglement of the total system state should be between 0 and 1, and the rank of the measurement should be greater than one to enable a disturbance of the (1,4) state by the second measurement. Additionally, we show that a zero $14|23$ bipartition entanglement resulting from a measurement that is an inseparable operator is not feasible. Our demonstration showcases the feasibility of generating entanglement between A and C for different measurement operators. Furthermore, an illustrative example underscores how successive measurements yield more substantial entanglement compared to a single measurement. Specifically, when employing one unentangled measurement incapable of establishing entanglement, a second measurement involving purely separable operation becomes effective. Our approach is not limited to bipartite qubit states but can also be generalized to higher-dimensional maximally entangled states.
@@ II GENERALIZED PROTOCOL
a pure state which is separable in $12|34$ bipartition. Now Bob’s first measurement $\mathcal{M}_{1}$ can be described by a set of two-qubit positive-semidefinite elements $\{\Pi_{n}\bigm{|}n=1,\ldots,K,\;\Pi_{n}\geq 0,\;\sum_{n}\Pi_{n}=\mathbb{I}_{4}\}$ where each POVM element $\Pi_{n}$ can be decomposed in the form
@@
where $|\phi_{n\alpha}\rangle$, $\alpha=1,2,3,4$ form a complete orthonormal basis on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $\pi_{n\alpha}\geq 0$ for $\alpha=1,2,3,4$. Here, $K$ represents the number of POVM elements in a POVM measurement. If Bob performs a joint two-qubit generalized quantum measurement $\mathcal{M}_{1}$ on ($2,3$) of the $|\Phi\rangle$ state then the post-measurement state of the $n$-th outcome is
@@
where $|\phi^{*}\rangle$ denotes the complex conjugate of $|\phi\rangle$ in the computational basis. Note that after the measurement, the total system state Eq.5 remains pure.
@@
$\rho_{14|n}$ $=$ $\frac{1}{4p_{n}}\sum_{\alpha=1}^{4}\pi_{n\alpha}|\phi_{n\alpha}^{*}\rangle\langle\phi_{n\alpha}^{*}|_{14}$ (6)
@@
where $\Pi_{n}^{*}=\sum_{\alpha=1}^{4}\pi_{n\alpha}|\phi_{n\alpha}^{*}\rangle\langle\phi_{n\alpha}^{*}|$. We also determine the post-measurement state of (1,2) and (3,4) pairs; for details, see Appendix A. Equation 6 illustrates how Bob’s measurement fully determines the state of the pair (1, 4). This equation also highlights the direct relationship between the post-measurement state of the (1,4) pair and the POVM measurement element. Conventional thought suggests that entanglement at the (1,4) pair requires an entangled measurement. However, we demonstrate that although an entangled POVM is sufficient to obtain an entangled state at (1,4), it is not a necessary condition
Our demonstration reveals that through a sequential approach, where Bob performs an initial measurement $\mathcal{M}_{1}$ and, if necessary, follows it with another measurement $\mathcal{M}_{2}$, entanglement can be achieved. This sequential strategy, coupled with appropriate post-processing after the first measurement, facilitates the establishment of entanglement between parties A and C. Notably, the second measurement $\mathcal{M}_{2}$ does not need to be an inseparable operation; a measurement that is separable operation can also accomplish the task.
@@
We consider the second measurement $\mathcal{M}_{2}$ is also two qubits generalized measurement (POVM) which can be specified by a collection of two qubits semi-definite operators $E_{m}$ satisfying $\sum_{m=1}E_{m}=\mathbb{I}_{4}$ where each POVM element $E_{m}$ can be decomposed in the previous form as
@@
Now using Eq.5, and Eq.7, the above Eq.8 can be written as
@@
where $|\phi_{n\alpha(\beta)}\rangle$, $\alpha(\beta)=1,2,3,4$ are orthonormal basis of the $n$-th element $\Pi_{n}$ of the first measurement $\mathcal{M}_{1}$.
@@
Next, to analyze the bipartition entanglement of post-measurement states for each outcome, we utilize the I-concurrence rungta for higher-dimensional pure quantum systems. Throughout this work, when referring to bipartition entanglement, we specifically mean the I-concurrence. The bipartition entanglement $A|B$ is defined using the following expression:
@@
Proof: A POVM, $\mathcal{M}$ is called as rank-1 when all its elements are rank-1 operators. Suppose $\mathcal{M}_{1}$ is a projective or rank-1 POVM, implying each element $\Pi_{n}$ of $\mathcal{M}_{1}$ measurement is rank-1, such that for $\alpha=1$, $\Pi_{n}=\pi_{n1}|\phi_{n1}\rangle\langle\phi_{n1}|$. Then, according to Eq.5, $|\Phi_{n}\rangle$ can be expressed as:
@@
Hence, if $\Pi_{n}$ is a projective or rank-1 entangled POVM element operator, the post-measurement state between A and C will be entangled, but no subsequent measurement can alter the entanglement in the (1,4) pair. Similarly, if $\Pi_{n}$ is a rank-1 unentangled POVM element or projective element, the post-measurement state between (1,4) will be separable, and no further sequential measurement on the (2,3) pair can entangle the (1,4) pair.
The above proof indicates that if Bob performs the first measurement with a rank greater than one, there is a possibility to disturb the (1,4) pair using another round of measurement on the (2,3) pair. This is because, from Eq.5, it can be understood that the $14|23$ bipartition may exhibit non-zero entanglement after Bob’s first measurement if its rank is greater than one. Consequently, one might pursue a subsequent measurement on the (2,3) pair to disturb the (1,4) pair.
Lemma-2: After performing the first measurement $\mathcal{M}_{1}$ on the (2,3) pair of the initial state Eq.1, if the bipartition entanglement $\mathcal{C}_{n|14vs23}$ of the four-qubit state $|\Phi_{n}\rangle$ (Eq.5) exhibits non-zero entanglement for a given outcome, i.e., if the $14|23$ bipartition entanglement lies within the range $0<\mathcal{C}_{n|14vs23}(|\Phi_{n}\rangle\langle\Phi_{n}|)\leq 1$, and the rank of the first measurement element is greater than one, then the second measurement $\mathcal{M}_{2}$ has the potential to disturb the entanglement of the (1,4) pair
@@
$=$ $\sqrt{\frac{4}{3}\left[1-\frac{\sum_{\alpha=1}^{4}\pi_{n\alpha}^{2}}{(\mbox{Tr}\Pi_{n})^{2}}\right]}$
@@
Initially, in the initial state (Eq.1) $12|34$ bipartition entanglement is zero i.e. $\mathcal{C}_{12vs34}(|\Phi\rangle\langle\Phi|)=0$ and in the $14|23$ bipartition $\mathcal{C}_{14vs23}(|\Phi\rangle\langle\Phi|)=1$. To get (1,4) pair entangled we need a measurement $\mathcal{M}_{1}$ that can produce non-zero entanglement in $12|34$ bipartition. So, until and unless Bob’s first measurement $\mathcal{M}_{1}$ produces entanglement in the $12|34$ bipartition we can not disturb the state between (1,4) pair during the second measurement even if $14|23$ bipartition has non-zero entanglement. Therefore, $\mathcal{C}_{n|12vs34}(|\Phi\rangle\langle\Phi|)>0$ is necessary for any type of measurement. However, this condition alone is not sufficient to entangle the (1,4) pair. There exist unentangled measurements capable of generating non-zero entanglement in the $12|34$ bipartition but unable to establish entanglement between qubits A and C. As previously established, if $\Pi_{n}$ is an unentangled element operator, the post-measurement state between the (1,4) pair will also be unentangled.
To get an entangled state between A and C, we have now given the protocols for the full two-qubit measurements of different measurement categories.
@@ II.1 Inseparable operations
An inseparable operation is defined as follows: let a quantum operation $\mathcal{E}_{A}$ acting on a system $\rho_{A}\in\mathcal{H}_{A}$ (Hilbert space), the Choi state is defined as choi :
$Choi(\mathcal{E}_{A})=(\mathbb{I}_{A_{1}}\otimes\mathcal{E}_{A})|\phi^{+}\rangle\langle\phi^{+}|_{A_{1}A}$ (15)
Here, $\mathbb{I}_{A_{1}}$ is the identity operator in $\mathcal{H}_{A_{1}}$ with dimension $\dim(\mathcal{H}_{A_{1}})=\dim(\mathcal{H}_{A})=d$. The state $|\phi^{+}\rangle$ is a normalized maximally entangled state in $\mathcal{H}_{A_{1}}\otimes\mathcal{H}_{A}$, given by $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle_{A_{1}}|i\rangle_{A}$ where $|i\rangle_{A_{1}}$ and $|i\rangle_{A}$ are orthonormal bases in $\mathcal{H}_{A_{1}}$ and $\mathcal{H}_{A}$, respectively.
Consider a quantum operation $\mathcal{M}_{AB}$ acting on a bipartite state $\rho_{AB}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. One can relabel the bases of $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ such that it becomes $\mathcal{H}^{d^{A}\times d^{B}}=\mathcal{H}_{C}$. Now, the whole operation can be viewed as a single-partite operation on $\mathcal{H}_{C}$, and one can represent it as $(\mathbb{I}_{C_{1}}\otimes\mathcal{M}_{C})|\phi^{+}\rangle\langle\phi^{+}|_{C_{1}C}$, where $|\phi^{+}\rangle_{C_{1}C}$ is the maximally entangled state in $\mathcal{H}_{C_{1}}\otimes\mathcal{H}_{C}$.
Alternatively, one can form two pairs of maximally entangled states of two qubits $|\phi^{+}\rangle\langle\phi^{+}|_{A_{1}A}$ and $|\phi^{+}\rangle\langle\phi^{+}|_{BB_{1}}$ and apply the quantum operation on their tensor product:
@@
If the entanglement of the state $\rho_{A_{1}ABB_{1}}$ across the bipartition $A_{1}A\ |\ BB_{1}$ is non-zero, then the operation is called an inseparable operation; otherwise, it is a separable operation cirac .
Both approaches are equivalent to define separable and inseparable quantum operations, as $|\phi^{+}\rangle\langle\phi^{+}|_{A_{1}A}\otimes|\phi^{+}\rangle\langle\phi^{+}|_{BB_{1}}=|\phi^{+}\rangle\langle\phi^{+}|_{CC_{1}}$. We follow the definition in Eq.16 to differentiate the separable and inseparable operations throughout the work.
In our work, we consider two initial states, $|\phi^{+}\rangle$, which are maximally entangled two-qubit states. The entanglement of the four-qubit post-measurement state, $|\Phi_{n}\rangle$, in the $12|34$ bipartition, i.e., $\mathcal{C}_{n|12vs34}(|\Phi_{n}\rangle\langle\Phi_{n}|)$, is the inseparability condition of the $\Pi_{n}$ element of any arbitrary two-qubit generalized measurement from an operational point of view. We find out the exact expression of inseparability criteria cirac for two-qubit arbitrary measurement using the I-concurrence relation rungta which is given by (for details see Appendix B):
@@
$=$ $\sqrt{\frac{4}{3}\left[1-\frac{1}{(\mbox{Tr}\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|ij}a^{*}_{n\eta|kl}a_{n\gamma|k^{\prime}l}\right]}.$